Last night, we put the IFO in FP Michelson configuration. We took transfer functions of CARM and DARM, first using CM excitations directly on the ETMs, and then using modulations of the laser frequency via MC excitation. We found that there was basically no coupling into DARM using the MC excitation, but that there was coherence in DARM using the ETM excitation. Therefore, I tuned the ETM common mode in the output matrix. I did this by taking transfer functions of PD1_Q with PD2_I (see attached plot). I changed the drdown_bang script to set C1:LSC-BTMTRX_14 0.98 and C1:LSC-BTMTRX_24 1.02.